# Program evaluation with remotely sensed outcomes

### v2

@@ 6 Recommendations for practice
Common empirical practice can be highly biased if the remotely sensed variable is post-outcome, i.e. caused by the outcome of interest. For example, changes in satellite images are caused by fires or variation in local income, but not vice versa. Theoretically, we demonstrate that the resulting bias can be positive or negative. Empirically, we demonstrate that common practice may attenuate the treatment effect in a real environmental application, and it may have positive or negative bias for the treatment effect in a semi-synthetic development application.
@@
However, the intuition underlying empirical work is powerful: the conditional distribution of the RSV given the outcome and treatment is stable across samples. We use such an assumption to nonparametrically identify treatment effects, and derive a novel formula. We demonstrate that it is empirically plausible in a real development application. Finally, we characterize the optimal way to use remotely sensed outcomes for program evaluation.
Our findings inform how to conduct program evaluation with RSVs in practice.
  • Auxiliary sample. Collect RSVs with linked outcomes in an auxiliary, observational sample. In this sample, the treatments may be missing or deterministic. It is helpful, but unnecessary, to collect some outcomes in the primary, experimental sample. The distribution of RSVs given outcomes and treatments should be stable across samples.
  • Three predictions. Use machine learning to predict not only the outcome, but also the treatment and sample indicator, given the RSV. Researchers can use complex machine learning methods with unknown statistical properties, while remaining agnostic about how the RSV is generated.
  • Robust inference. Use Algorithm 4.1 to aggregate the three predictions into an efficient representation of the RSV for program evaluation. Algorithm 4.1 is robust to mis-specification; it provides valid inference as long as the learned representation converges at any rate to any limit that is correlated with outcome variation.
Our framework clarifies the key diagnostics that empirical researchers should assess.
  • RSV relevance test. Evaluate whether the learned RSV representation is correlated with outcome variation. Tests for weak instruments can detect weak RSVs.
  • Stability and no direct effect joint test. If the treatment effects estimated with two different RSV representations are significantly different, then reject the null hypothesis that both conditions hold: stability (Assumption 2.2), and no direct effects (Assumption 2.3(ii)).
  • RSV density plots. If possible, collect some experimental outcomes. If the conditional density of the RSV given the outcome and treatment is similar across experimental and observational units, then stability (Assumption 2.2) is plausible. If it is similar across treated and untreated experimental units, then no direct effects (Assumption 2.3(ii)) is plausible.
In summary, researchers may substantially reduce research costs by using post-outcome RSVs instead of directly measured outcomes, without compromising valid inference.
@@
Our framework poses new questions, e.g. how to jointly design experiments (treatment assignment), surveys (outcome collection), and sensors (RSV extraction) to optimally estimate treatment effects, subject to cost constraints. Our main result (Theorem 3.1) may inform not only the use of satellite images, but also the design of cheap, noisy surveys. Future work may also extend our inference guarantees (Propositions 4.1 and 4.2) to accommodate high dimensional outcomes.
As remote sensing expands the frontier of data availability in economics, our framework provides practical principles for its use in program evaluation.
Appendix A Related work and model details
@@
Figure 8: Remotely sensed variables are increasingly popular in published papers.
@@
Figure 8 illustrates the increasing popularity of RSVs in empirical research. First, we collected papers published in the AEA Journals, Econometrica, Quarterly Journal of Economics, Review of Economic Studies and Journal of Political Economy using a keyword search of “remotely sensed variables”, “mobile phone”, “satellite”, “machine learning”, and “drones / aerial” on their websites. Next, we collected papers published in Nature and Science using a keyword search of “remotely sensed variables” on their websites. We subsetted to the papers with RSVs in their main empirical analysis.
$D$$R$$Y$
Figure 9: Causal graph of surrogacy model.
@@
Figure 9 illustrate the causal graph associated with the surrogacy identifying assumptions (Prentice,, 1989; Athey et al.,, 2024): the surrogate $R$ fully mediates the effect of the treatment $D$ on the outcome $Y$. Our RSV identifying assumptions are the opposite, as illustrated by Figure 3.
Table 3: Implications of RSV identifying assumptions.
@@
2.1 $\Pr\{Y(d)|S,X,D\}$ $\Pr\{Y(d)|S=e,X\}$ $\Pr\{Y(d)|S=o,X,D\}$ Differs across samples
2.2 $\Pr(R|S,X,D,Y)$ $\Pr(R|X,D,Y)$ $\Pr(R|X,D,Y)$ Stable across samples
2.3(i) $\Pr(R|S,X,D,Y)$ $\Pr(R|X,D,Y)$ $\Pr(R|X,D,Y)$ Differs across treatments if complete cases
2.3(ii) $\Pr(R|S,X,D,Y)$ $\Pr(R|X,Y)$ $\Pr(R|X,Y)$ Stable across treatments if incomplete cases
Table 3 summarizes the implications of our identifying assumptions. Our identifying assumptions allow the propensity score $\Pr(D=1\mid S,X)$ to differ across samples.
Appendix B Crop burning illustration
B.1 Characterizing bias with binary variables
@@
Suppose Assumptions 2.12.2, and 2.3(ii) hold, as well as Assumption B.1. Then, without loss of generality, $\mathbb{E}(Y|R,S=o)=\tilde{\beta}_{0}+\tilde{\beta}R$, $\mathbb{E}(R|Y,D,S=e)=\beta_{0}+\beta Y$, and $\theta_{0}=\mathbb{E}(Y|D=1,S=e)-\mathbb{E}(Y|D=0,S=e)$ for some scalars $(\tilde{\beta}_{0},\tilde{\beta},\beta_{0},\beta)$.
@@
$=\tilde{\beta}\beta\{\mathbb{E}(Y|D=1,S=e)-\mathbb{E}(Y|D=0,S=e)\}=\tilde{\beta}\beta\theta_{0}.\qed$
@@
Proposition B.2 (Binary).
@@
$\tilde{\beta}\{\mathbb{E}(R|D=1,S=e)-\mathbb{E}(R|D=0,S=e)\}=\tilde{\beta}\beta\theta_{0}.\qed$
Within the main text, we take $\tilde{\beta}=1$ so that $\widetilde{\theta}=\beta\theta_{0}$ by Corollary B.1 since Jack et al., (2025) directly plug-in the RSV $R$ as the outcome in the experiment.
B.2 Implementation details
We present details for the empirical example in Section 3.2, based on Jack et al., (2025). We classify a field as experimental ($S=e$) if it did not receive a random spot check. We classify a field as observational ($S=o$) if it received a random spot check. This example has complete cases (Assumption 2.3(ii)); the treatment varies in the observational sample.
In Section 3.2, we define $R\in\{0,1\}$ as the authors’ “maximum accuracy” classifier for whether a field has not been burned. Jack et al., (2025) construct another “balanced accuracy” classifier based on an alternative threshold rule. We find similar results in Table 4, now defining $R\in\{0,1\}$ as the authors’ “balanced accuracy” classifier.
Table 4: Underestimation of treatment effects in crop burning experiment: Revisited.
Common practice Bias Causal parameter
Estimand $\widetilde{\theta}$ $\beta$ $\theta$
Estimate 0.074 0.601*** 0.123
(0.049) (0.068) (0.086)
Notes: The RSV is the field-level “balanced accuracy” label defined by Jack et al., (2025), which applies a threshold rule to the predicted probability of not being burned. The “observational sample” has fields that received a random spot check, and the “experimental sample” has other fields. For illustration, we conduct linear estimation, controlling for stratum fixed effects. Standard errors are based on $5000$ bootstrap replications clustered at the village level.
(a) Experimental units only.
(b) Auxiliary sample: Observational units.
Figure 10: We illustrate the samples in our re-analysis of the crop burning experiment in Jack et al., (2025), plotting the map of villages in Bathinda and Faridkot, which are two districts in Punjab, India. Experimental villages are those in which less than 50% of fields received a spot check. Observational villages are those in which more than 50% of fields received a spot check.
Appendix C Identification proofs
C.1 Proof of Proposition 3.1
@@ Appendix C Identification proofs
$=\int\Pr(Y=1\mid R=r,D=1,S=e)\Pr(R=r\mid D=1,S=e)\mathrm{d}r,$
where (1) follows by Assumption 2.1. Next we rewrite the implicit target
$\widetilde{\mu}(1):=\int\Pr(Y=1\mid R=r,S=o)\Pr(R=r\mid D=1,S=e)\mathrm{d}r.$
In particular,
$\Pr(Y=1\mid R,S=o)$ =$\overset{(1)}{=}\frac{\Pr(R\mid Y=1,S=o)\Pr(Y=1\mid S=o)}{\Pr(R\mid S=o)}$
=$\overset{(2)}{=}\frac{\Pr(R\mid Y=1,D=1,S=e)P(Y=1\mid S=o)}{\Pr(R\mid S=o)}$
=$\overset{(3)}{=}\Pr(Y=1\mid R,D=1,S=e)\frac{\Pr(R\mid D=1,S=e)\Pr(Y=1\mid S=o)}{\Pr(Y=1\mid D=1,S=e)\Pr(R\mid S=o)}$
=$\overset{(4)}{=}\Pr(Y=1\mid R,D=1,S=e)\frac{\Pr(Y=1\mid S=o)}{\Pr\{Y(1)=1\mid S=e\}}\frac{\Pr(R\mid D=1,S=e)}{\Pr(R\mid S=o)}$
@@
$\frac{\Pr\{Y(0)=1\mid S=e\}}{\Pr\{Y(1)=1\mid S=e\}}\int\frac{\Pr(R=r\mid D=1)}{\Pr(R=r\mid D=0)}\Pr(Y=1\mid R=r,D=1,S=e)\Pr(R=r\mid D=1,S=e)\mathrm{d}r,$
and $\widetilde{\mu}(1)-\mu(1)$ equals
$\int\left[\frac{\Pr\{Y(0)=1\mid S=e\}}{\Pr\{Y(1)=1\mid S=e\}}\frac{\Pr(R=r\mid D=1)}{\Pr(R=r\mid D=0)}-1\right]\Pr(Y=1\mid R=r,D=1,S=e)\Pr(R=r\mid D=1,S=e)\mathrm{d}r.$
Further rearranging, we can write
$\Pr(Y=1\mid R,D=1,S=e)\Pr(R\mid D=1,S=e)$ $=\Pr(R\mid Y=1,D=1,S=e)\Pr\{Y(1)=1\mid S=e\}.$
@@
$\mu(1)\int\left[\frac{\Pr\{Y(0)=1\mid S=e\}}{\Pr\{Y(1)=1\mid S=e\}}\frac{\Pr(R=r\mid D=1)}{\Pr(R=r\mid D=0)}-1\right]\Pr(R=r\mid Y=1,D=1,S=e)\mathrm{d}r.$
We can follow a similar argument for $\widetilde{\mu}(0)$. In particular, $\widetilde{\mu}(0)$ equals
$\int\Pr(Y=1\mid R=r,D=0,S=e)\frac{\Pr(Y=1\mid S=o)}{\Pr\{Y(0)=1\mid S=e\}}\frac{\Pr(R=r\mid D=0,S=e)}{\Pr(R=r\mid S=o)}\Pr(R=r\mid D=0,S=e)\mathrm{d}r$
$=\int\Pr(Y=1\mid R=r,D=0,S=e)\Pr(R=r\mid D=0,S=e)\mathrm{d}r=\mu(0).$
@@
$\Pr\{Y(0)$ $=1\mid S=e\}=\Pr\{Y(0)=1\mid D=0,S=e\}=\Pr(Y=1\mid D=0,S=e)$
$=\Pr(Y=1\mid D=0,S=o)=\Pr(Y=1\mid S=o),$
as well as $\Pr(R=r\mid D=0,S=e)=\Pr(R=r\mid D=0)$ as before.
@@
$R\mid Y,D,S=\begin{cases}Y\mbox{ with probability }1/2\\ 0\mbox{ otherwise}.\end{cases}$
@@
$\Pr(R=1\mid D=1)=\int\Pr(R=1,Y=y\mid D=1,S=e)\mathrm{d}y$
$=\int\Pr(R=1\mid Y=y,D=1,S=e)\Pr(Y=y\mid D=1,S=e)\mathrm{d}y$
$=\Pr(Y=1\mid D=1,S=e)+\frac{1}{2}\Pr(Y=0\mid D=1,S=e)$
$=\Pr\{Y(1)=1\mid S=e\}+\frac{1}{2}\Pr\{Y(1)=0\mid S=e\}=\frac{1}{2}[1+\Pr\{Y(1)=1\mid S=e\}]$

### v3

@@ 6 Recommendations for Practice
Common empirical practice can be highly biased when the remotely sensed variable is post-outcome, i.e., when the RSV is caused by the outcome of interest. For example, changes in satellite images are caused by fires or variation in local income, but not vice versa. Theoretically, we demonstrate that this bias can have arbitrary sign. Empirically, we find that common practice may attenuate the treatment effect in a real environmental application, and it may have positive or negative bias in a semi-synthetic development application.
@@
However, the core intuition underlying empirical work is powerful: the conditional distribution of the RSV given the outcome and treatment should be stable across samples. By formalizing this intuition, we develop a framework that nonparametrically identifies treatment effects and efficiently uses RSVs for program evaluation.
We conclude with some concrete recommendations for researchers conducting program evaluation with remotely sensed outcomes. We organize our recommendations into (i) implementation steps and (ii) diagnostic tests.
6.1 Three Steps for Robust Inference with RSVs
To apply our framework, researchers should follow three steps when designing studies with remotely sensed outcomes.
Step 1: Construct a credible observational sample. Build an auxiliary, observational sample that contains RSVs with linked outcomes. The key identifying assumption is stability: the conditional distribution of the RSV given the covariates, treatment, and outcome should be stable across the experimental and observational samples. Consequently, the researcher’s most important design decision is how to construct an observational sample in a way that makes the stability assumption defensible.
Stability is most credible if the researcher randomly selects units for the experimental sample and observational sample prior to conducting the experiment. When random selection is infeasible, researchers should construct an observational sample that is similar to the experimental sample along a few dimensions. (i) Select units that are geographically close to the experimental sample. (ii) Select units with similar baseline covariates, e.g. rural-urban compositions, to the experimental sample. (iii) Select units that align with the timing of the experimental sample. That is, in the observational sample, collect outcomes measured in the same period as the experiment’s endline, and collect RSVs measured at the same time that RSVs will be collected in the experiment. If exact alignment is not possible, then match the time difference so that the temporal relationship might be preserved.
Step 2: Estimate three predictions from RSVs. Efficient use of RSVs for downstream causal inference requires three predictions: the outcome, treatment, and sample indicator given the RSV. Common practice uses only the first prediction, in a way that sacrifices not only accuracy but also precision. In other words, by using all three predictions, our method improves not only bias but also variance. Researchers can construct these predictions using complex deep learning algorithms with unknown statistical properties.
Step 3: Apply robust inference algorithm. Algorithm 1 aggregates the three predictions into an efficient scalar representation of the high-dimensional RSV (where efficiency is within the class of models satisfying the conditional moments we derive). This aggregation reduces the dimension while preserving information relevant for causal inference. Critically, Algorithm 1 is robust to misspecification; it provides valid, $n^{-1/2}$ inference even if all three predictions are misspecified. It only requires that the learned representation converges at any rate to any limit that predicts outcome variation. This robustness property justifies the use of deep neural networks, which are widely implemented by empirical researchers in this literature, while providing the safeguard of a principled inferential framework.
6.2 Three Diagnostics for RSVs
In addition to the point estimate and confidence interval of Algorithm 1, researchers should report the following diagnostics to defend the credibility of their RSV.
Diagnostic 1: RSV relevance test. Just as weak instruments threaten inference in instrumental variable models, weak RSVs—those that poorly predict the outcome—threaten inference in our framework. Researchers should assess whether the correlation between their estimated representation and the outcome is significantly different than zero. Figure 6 illustrates this diagnostic, showing that satellite images are strongly related to village-level poverty. When relevance is weak, researchers should either improve RSV measurements (e.g., by using higher resolution images or additional data sources) or acknowledge that inference is inconclusive.
Diagnostic 2: Joint test of identifying assumptions. Our identifying assumptions are jointly testable. If two different RSV representations yield significantly different treatment effect estimates, then reject the null hypothesis that our identifying assumptions jointly hold: stability (Assumption 2) and no direct effects (Assumption 3(ii)). This is essentially a specification test. Researchers should verify that their results are robust to alternative ways of representing the RSV, such as using different machine learning architectures or different subsets of RSV features.
Diagnostic 3: RSV density plots if possible. When outcomes are observed for some experimental units, researchers can directly visualize whether stability holds by comparing conditional densities. Specifically, plot the RSV density in the experimental sample against the RSV density in the observational sample, for each combination of treatment and outcome values. If these densities align, then stability is plausible, as in Figure 2. Similarly, researchers can assess the no-direct-effects assumption (Assumption 3(ii)) by comparing the RSV density of treated units against the RSV density of untreated units, for each outcome value. If treatment only affects the RSV via the outcome, these densities should coincide, as in Figures H.1 and H.3.
6.3 Broader Implications
By following these steps, researchers can conduct rigorous program evaluation while substantially reducing costs. In our semi-synthetic exercise, we recover treatment effects and confidence intervals comparable to those obtained with direct outcome measurements, despite using satellite images for half the sample.
Our framework generalizes beyond simple randomized experiments. Appendix J extends our identification results to quasi-experiments (e.g., instrumental variables and difference-in-differences). The key insight remains the same: stability in how the RSV relates to other variables across samples allows us to transport this relationship to the sample where the outcome is missing.
@@ 6.3 Broader Implications
Our framework poses new questions. Future research may study a new experimental design problem: how to simultaneously design treatment assignment, outcome collection, and sensor deployment to maximize statistical power subject to budget constraints. Our main identification result provides the foundation to set up this problem. Our main result may inform not only the use of satellite images, but also the design of cheap, noisy surveys. Finally, future work may extend our inference guarantees to accommodate high-dimensional outcomes.
As remote sensing expands the frontier of data availability in economics, our framework offers practical principles for its use in program evaluation. This paper corrects a widespread but biased practice and provides concrete recommendations on how to incorporate satellite images and other remotely sensed data into causal inference. These new data sources and new methods broaden access to rigorous impact evaluation in cost-constrained environments.
Appendix Materials
Ashesh Rambachan, Rahul Singh, Davide Viviano
Appendix A provides additional figures and tables that motivate our model. Appendix B derives the bias of common practice. Appendices C and D prove our main identification and inference results, respectively. Appendices E and F extend these results from binary outcomes to discrete and continuous outcomes, respectively. Appendix G provides details for the crop burning illustration. Appendices H and I provide details and extensions, respectively, for the Smartcard illustration. Finally, Appendix J extends our identification framework to quasi-experimental settings.
Appendix A Related Work and Additional Model Details
@@
Figure A.1: Remotely sensed variables are increasingly popular in published papers.
@@
Figure A.1 illustrates the increasing popularity of RSVs in empirical research. First, we collected papers published in the AEA journals, Econometrica, Quarterly Journal of Economics, Review of Economic Studies and Journal of Political Economy using a keyword search of “remotely sensed variables”, “mobile phone”, “satellite”, “machine learning”, and “drones / aerial” on their websites. Next, we collected papers published in Nature and Science using a keyword search of “remotely sensed variables” on their websites. We subsetted to the papers with RSVs in their main empirical analysis.
DDRRYY
Figure A.2: Causal graph of surrogacy model.
@@
Figure A.2 illustrate the causal graph associated with the surrogacy identifying assumptions (Prentice, 1989; Athey et al., 2025): the surrogate $R$ fully mediates the effect of the treatment $D$ on the outcome $Y$. Our RSV identifying assumptions are the opposite, as illustrated by Figure 3.
Table A.1: Implications of RSV identifying assumptions.
@@
1 $\Pr\{Y(d)|S,X,D\}$ $\Pr\{Y(d)|S=e,X\}$ $\Pr\{Y(d)|S=o,X,D\}$ Differs across samples
2 $\Pr(R|S,X,D,Y)$ $\Pr(R|X,D,Y)$ $\Pr(R|X,D,Y)$ Stable across samples
3(i) $\Pr(R|S,X,D,Y)$ $\Pr(R|X,D,Y)$ $\Pr(R|X,D,Y)$ Differs across treatments if complete cases
3(ii) $\Pr(R|S,X,D,Y)$ $\Pr(R|X,Y)$ $\Pr(R|X,Y)$ Stable across treatments if incomplete cases
Table A.1 summarizes the implications of our identifying assumptions. Our identifying assumptions allow the propensity score $\Pr(D=1\mid S,X)$ to differ across samples.
Appendix B Bias of Common Practice Proofs
B.1 Bias with Binary Variables
@@
Suppose Assumptions 12, and 3(ii) hold, as well as Assumption B.1. Then, without loss of generality, $\mathbb{E}(Y|R,S=o)=\tilde{\beta}_{0}+\tilde{\beta}R$, $\mathbb{E}(R|Y,D,S=e)=\beta_{0}+\beta Y$, and $\theta_{0}=\mathbb{E}(Y|D=1,S=e)-\mathbb{E}(Y|D=0,S=e)$ for some scalars $(\tilde{\beta}_{0},\tilde{\beta},\beta_{0},\beta)$.
@@
$=\tilde{\beta}\beta\{\mathbb{E}(Y|D=1,S=e)-\mathbb{E}(Y|D=0,S=e)\}=\tilde{\beta}\beta\theta_{0}.$
@@
Proposition B.2 (Identification with binary variables).
@@
$\tilde{\beta}\{\mathbb{E}(R|D=1,S=e)-\mathbb{E}(R|D=0,S=e)\}=\tilde{\beta}\beta\theta_{0}.$
B.2 Bias with Random Sample Selection
As in Proposition 1, consider the setting with binary outcomes, no covariates, and incomplete cases. We now verify that the common practice is biased, even when the researcher randomly assigns units to either the experimental or the observational sample.
In this thought experiment, the researcher implements the following study: (i) randomization of units into the experimental or observational sample; (ii) randomization of experimental units to treatment or control; (iii) the common practice described in Section 3.1. No observational units are treated.
To prove this claim, it is helpful to explicitly write the potential outcome as $Y(d)$ and the potential RSV as $R(y)$. The potential RSV is not indexed by the treatment under Assumption 3(ii). Each unit is characterized by the random vector $\{S,D,Y(0),Y(1),R(0),R(1)\}.$
Proposition B.3 (Bias of common practice with random sample selection).
Suppose Assumptions 1, 2, and 3(ii) hold with $X=\varnothing$. Further assume
  1. i.
    SUTVA for RSV: R=YR(1)+(1Y)R(0)R=YR(1)+(1-Y)R(0) almost surely.
  2. ii.
[…]
  3. iii.
[…]
  4. iv.
    Observational sample is untreated: $\Pr(D=0|S=o)=1$.
Despite randomized sample selection, the bias of common practice persists. In particular, the conclusions of Proposition 1 continue to hold.
Proof.
[…]
  1. 1.
[…]
  2. 2.
[…]
    $\mathbb{E}\{h(Y,R)|D=0,S=e\}$ $=\mathbb{E}[h\{Y,YR(1)+(1-Y)R(0)\}|D=0,S=e]$
    $=\mathbb{E}(h[Y(0),Y(0)R(1)+\{1-Y(0)\}R(0)]|D=0,S=e)$
    $=\mathbb{E}(h[Y(0),Y(0)R(1)+\{1-Y(0)\}R(0)]|S=e)$
    $=\mathbb{E}(h[Y(0),Y(0)R(1)+\{1-Y(0)\}R(0)]).$
    By SUTVA, untreated observational units, and randomized sample selection,
    $\mathbb{E}\{h(Y,R)|D=0,S=o\}$ $=\mathbb{E}[h\{Y,YR(1)+(1-Y)R(0)\}|D=0,S=o]$
    $=\mathbb{E}(h[Y(0),Y(0)R(1)+\{1-Y(0)\}R(0)]|D=0,S=o)$
    $=\mathbb{E}(h[Y(0),Y(0)R(1)+\{1-Y(0)\}R(0)]|S=o)$
    $=\mathbb{E}(h[Y(0),Y(0)R(1)+\{1-Y(0)\}R(0)]).$
B.3 Proof of Proposition 1
We characterize the bias. For simplicity, we take $X=\varnothing$. We allow $R$ to be discrete or continuous, and impose that $Y$ is binary for the sake of exposition. When constructing adversarial data-generating processes, we will choose simple examples with binary $R$, again for the sake of exposition. We proceed in steps.
  1. 1.
@@ B.3 Proof of Proposition 1
    $=\int\Pr(Y=1\mid R=r,D=1,S=e)f_{R}(r\mid D=1,S=e)\mathrm{d}r,$
    where (1) follows by Assumption 1.
  2. 2.
    Next we rewrite the implicit target
    $\widetilde{\mu}(1):=\int\Pr(Y=1\mid R=r,S=o)f_{R}(r\mid D=1,S=e)\mathrm{d}r.$
    Within this expression,
    $\Pr(Y=1\mid R,S=o)$ $\overset{(1)}{=}\frac{f_{R}(R\mid Y=1,S=o)\Pr(Y=1\mid S=o)}{f_{R}(R\mid S=o)}$
    $\overset{(2)}{=}\frac{f_{R}(R\mid Y=1,D=1,S=e)P(Y=1\mid S=o)}{f_{R}(R\mid S=o)}$
    $\overset{(3)}{=}\Pr(Y=1\mid R,D=1,S=e)\frac{f_{R}(R\mid D=1,S=e)\Pr(Y=1\mid S=o)}{\Pr(Y=1\mid D=1,S=e)f_{R}(R\mid S=o)}$
    $\overset{(4)}{=}\Pr(Y=1\mid R,D=1,S=e)\frac{\Pr(Y=1\mid S=o)}{\Pr\{Y(1)=1\mid S=e\}}\frac{f_{R}(R\mid D=1,S=e)}{f_{R}(R\mid S=o)}$
@@
    $\frac{\Pr\{Y(0)=1\mid S=e\}}{\Pr\{Y(1)=1\mid S=e\}}\int\frac{f_{R}(r\mid D=1)}{f_{R}(r\mid D=0)}\Pr(Y=1\mid R=r,D=1,S=e)f_{R}(r\mid D=1,S=e)\mathrm{d}r.$
  3. 3.
    Combining results, the bias for the treated potential outcome, $\widetilde{\mu}(1)-\mu(1)$, equals
    $\int\left[\frac{\Pr\{Y(0)=1\mid S=e\}}{\Pr\{Y(1)=1\mid S=e\}}\frac{f_{R}(r\mid D=1)}{f_{R}(r\mid D=0)}-1\right]\Pr(Y=1\mid R=r,D=1,S=e)f_{R}(r\mid D=1,S=e)\mathrm{d}r.$
    Within this expression,
    $\Pr(Y=1\mid R,D=1,S=e)f_{R}(R\mid D=1,S=e)=f_{Y,R}(Y=1,R\mid D=1,S=e)$
    $=f_{R}(R\mid Y=1,D=1,S=e)\Pr(Y=1\mid D=1,S=e\}$
    $=f_{R}(R\mid Y=1,D=1,S=e)\Pr\{Y(1)=1\mid S=e\}$
    $=f_{R}(R\mid Y=1,D=1,S=e)\mu(1).$
    by the law of total probability and Assumption 1.
@@
    $\mu(1)\int\left[\frac{\Pr\{Y(0)=1\mid S=e\}}{\Pr\{Y(1)=1\mid S=e\}}\frac{f_{R}(r\mid D=1)}{f_{R}(r\mid D=0)}-1\right]f_{R}(r\mid Y=1,D=1,S=e)\mathrm{d}r.$
@@
  4. 4.
    We can follow a similar argument for $\widetilde{\mu}(0)$. As in Step 2,
    $\widetilde{\mu}(0)=\int\Pr(Y=1\mid R=r,S=o)f_{R}(r\mid D=0,S=e)\mathrm{d}r$
    $=\int\Pr(Y=1\mid R=r,D=0,S=e)\frac{\Pr(Y=1\mid S=o)}{\Pr\{Y(0)=1\mid S=e\}}\frac{f_{R}(r\mid D=0,S=e)}{f_{R}(r\mid S=o)}f_{R}(r\mid D=0,S=e)\mathrm{d}r$
[…]
    $\widetilde{\mu}(0)=\int\Pr(Y=1\mid R=r,D=0,S=e)f_{R}(r\mid D=0,S=e)\mathrm{d}r=\mu(0)$
    where the second equality uses the same argument as Step 1.
  5. 5.
@@
    $R\mid Y,D,S=\begin{cases}Y\mbox{ with probability }1/2\\ 1\mbox{ otherwise}.\end{cases}$
@@
    $\Pr(R=1\mid D=d)$ $=\Pr(R=1\mid D=d,S=e)$
    $=\int\Pr(R=1,Y=y\mid D=d,S=e)\mathrm{d}y$
    $=\int\Pr(R=1\mid Y=y,D=d,S=e)\Pr(Y=y\mid D=d,S=e)\mathrm{d}y$
    $=\Pr(Y=1\mid D=d,S=e)+\frac{1}{2}\Pr(Y=0\mid D=d,S=e)$
    $=\Pr\{Y(d)=1\mid S=e\}+\frac{1}{2}\Pr\{Y(d)=0\mid S=e\}$
    $=\frac{1}{2}[1+\Pr\{Y(d)=1\mid S=e\}]$